mark can remember only the first 3 digits of his friends phone number, he also know that the number has 7 digits and that the last digit is not a 0. If mark were to dial all of the possible numbers and if it takes him 23 seconds to try each one, how long would it take it to try every possibility?
Answer by sudhanshu_kmr(1152)


here we have to find total numbers of 4 digits, where 0 is not on last.

total no. of possible numbers = 10 * 10 *10 *9 = 9000


time= 9000 * 23 seconds